120g of Na2CO3 Salt was added to 80cm3 of water at 30oC. The solubility of the 30oC is 8.0mol/dm3. Calculate the mass of the salt which remained un dissolved. (Na=23; C=12; O=16)
1
Expert's answer
2020-06-30T07:47:25-0400

1.find the n of the Na2CO3:

"n=\frac{m}{M}=\frac{120}{106}=1.13"

2.find the amount of salt in the solution at 30 degrees

"n=C\times V=8*0.08=0.64"

3.find the amount of salt undissolved at 30 degrees

1.13-0.64=0.49

4.find the mass of salt undissolved at 30 degrees

"m=n\times M=0.49\times106=51.94"